OpenStudy (anonymous):

What is the exact area of a circle with a radius of m? A. 361pi m² B. 57pi m² C. 38pi m² D. 19pi m²

OpenStudy (anonymous):

a radius of?

OpenStudy (anonymous):

What is the value of the radius?

OpenStudy (anonymous):

sorry the radius is 19

OpenStudy (anonymous):

formula for the circle=pi(19)^2

OpenStudy (anonymous):

are those the answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Use the formula just that don't multiple the pi.Just the radius!

OpenStudy (anonymous):

because in the answer choices theres pi

OpenStudy (anonymous):

You just want to multiply the r^2 where r=radius

OpenStudy (anonymous):

so wat the answer

OpenStudy (anonymous):

lol you have to do it haha

OpenStudy (anonymous):

Omg just multiply Sir

OpenStudy (anonymous):

We have basically shown you all the work

OpenStudy (anonymous):

common I have to go to sleep

OpenStudy (anonymous):

multiply 19^2 and that is your answer

OpenStudy (anonymous):

19^2 = 19*19

OpenStudy (anonymous):

lol I do too :)

OpenStudy (anonymous):

thanks peop;e

OpenStudy (anonymous):

did you get your answer?